# Interpreting Key Mechanisms of Factual Recall in Transformer-Based Language Models

### v1

@@ 3.2 Attention Heads Pass Arguments to the Implicit Function
Figure 3: (a) The probability of $X$ and $Y$ when early unembed each layer’s outputs. The shaded regions indicate the standard deviation. (b) The effect of patching path $\textbf{a}^{l,h}_{t=-1}\rightarrow\textbf{r}^{11}_{\text{post}}$. (c) Value weighted attention pattern of L9H8 and L10H0. (d) For L9H8 and L10H0, their attention to $X$ is proportional to the projection value of their output along $\textbf{W}_{U}[X]$, indicating they are moving country names to the final position.
@@
We replicated experiments conducted by Merullo et al. (2023) using GPT-2 small. Given an input $p_{\text{original}}$, we utilize the $\texttt{Unembed}(\cdot)$ function to unembed each layer’s outputs $\textbf{r}^{l}_{\text{post},t=-1}$. Then, we apply softmax to the early-unembedded logits to assess the probabilities of $X$ and $Y$, respectively. Figure 3(a) illustrates the probability dynamics across layers. A significant change occurs at layer 9, where $X$ becomes noticeable in the residual stream. This marks the argument formation phase. By the end of layer 10, the probability of $Y$ surpasses that of $X$ for the first time, indicating the function application. Meanwhile, we implement the “path patching” algorithm, an activation patching algorithm introduced by Wang et al. (2023) to identify influential nodes. For patching, $p_{\text{original}}$ is sampled from the test dataset, while the set of interventions $\{p_{\text{intervention},n\leq N}\}$ is the entire test dataset. We employ Eq.6 as our impact metric. Path patching results illustrated in Figure 3(b) indicate that paths from outputs of L9H8 and L10H0, i.e., $\textbf{a}^{9,8}_{t=-1}$ and $\textbf{a}^{10,0}_{t=-1}$, to $\textbf{r}^{11}_{\text{post},t=-1}$ have the most substantial positive impact on correct prediction. The alignment between these experiments suggests that L9H8 and L10H0 play pivotal roles in the two crucial phases.
We identified that L9H8 and L10H0 are “mover heads”, which have garnered attention in many studies (Wang et al., 2023; Lieberum et al., 2023; Merullo et al., 2024). These heads are characterized by their tendency to move the information from any position they focus on to the final position (i.e., $t=-1$). As illustrated in Figure 3(c), both L9H8 and L10H0 predominantly attend to the token $X$; In Figure 3(d), we present a scatter plot illustrating the attention score from the final token to $X$, plotted against the inner product $<\textbf{a}^{l,h}_{t=-1},\textbf{W}_{U}[X]>$. The significant correlation between the attention score and the inner product (with Pearson correlation coefficient $0.77$ for L9H8 and $0.90$ for L10H0) strongly suggests that these two heads are responsible for passing the country name, the “argument” in the country-capital task, to the final position.
These pieces of evidence imply that the behavior exhibited by L9H8 and L10H0 can be likened to passing arguments to an implicit function, for example, “get_capital(x).” Moreover, beyond the mere act of argument passing, we found that L9H8 demonstrates some capacity to directly produce the correct answer, as discussed below.
@@
To examine L9H8 closely, we employ $\texttt{Unembed}(\cdot)$ to early-decode $(\textbf{W}^{\text{L9H8}}_{OV}\cdot\textbf{r}^{9}_{\text{pre},t=X})$ to obtain top-5 tokens. We noticed that correct capital cities emerged among these top tokens. This indicates that partial function application happens within this attention head. Mapping countries to capitals is an inherent ability of L9H8: Despite receiving the embedding vector of a country rather than a representation from deeper layers at position $X$, its OV matrix still yields some capital cities. Our investigation reveals that L9H8 is a task-specific head mainly activated given country-related prompts, as evidenced by its inactivation in other tasks (see appendix D). In contrast, although L10H0 lacks an inherent ability to map countries to capital cities within its OV matrix, it is activated across a variety of tasks, even those solely reliant on in-context learning (Wang et al., 2023; Merullo et al., 2024). It functions more like an amplifier for L9H8 or, more generally, an amplifier for the confident answer in the residual stream.
@@ 3.3 MLP Acts as the “Activation” of Attention Heads
In §3.2, we have demonstrated that only a few attention heads have a positively direct effect to the correct prediction. This naturally leads to a question: Given that Transformer models aggregate the outputs of all attention heads with equal weight (as shown in Eq. 1), how does the model discern and extract information written by specific heads within the residual stream? We first compare the norm of each attention head’s outputs and employ principal component analysis (PCA) to detect anomalous heads’ outputs. Unfortunately, these experiments failed to yield any informative results. Then, we hypothesize the existence of an “activation” mechanism after these attention heads that either amplifies or suppresses the outputs of certain attention heads. Now, let us turn to the subsequent module following the attention heads—the MLP layer.
Recall that a Transformer layer’s output $\textbf{r}^{l}_{\text{post}}$ is the linear combination of $\textbf{r}^{l}_{\text{pre}}$, the collective output of attention heads $\sum^{H-1}_{h=0}\textbf{a}^{l,h}$, and the MLP output $\textbf{m}^{l}$ (See Eq.1). This prompts us to consider whether MLP outputs can be decomposed into a linear combination of $\textbf{r}^{l}_{\text{pre}}$ and $\sum^{H-1}_{h=0}\textbf{a}^{l,h}$, where each term’s coefficient indicates the degree of amplification or erasure imposed by the MLP? Intrigued by this notion, we tackle the following multivariable linear regression problem:
@@
Here, $w^{l,\cdot}$ are scalars, and the intercept $\textbf{b}^{l}\in\mathbb{R}^{d_{\text{model}}}$ is expected to encapsulate higher-order transformations of head outputs, along with abstract information such as knowledge introduced by the MLP. We employ the gradient descent method to solve this linear regression problem, with the objective of minimizing the Mean Squared Error (MSE) loss between $\tilde{\textbf{m}}^{l}$ and $\textbf{m}^{l}$. We set the learning rate to 0.005 and performed 60,000 optimization steps using the SGD optimizer with a momentum of 0.99. Note that Eq.7 is position-specific, and our primary focus is on the final position in this section. Unless otherwise specified, the position subscripts of vectors default to $t=-1$ throughout the subsequent paper.
We employed 4-fold cross-validation to validate the regression solution. We found that, for each layer, the optimization consistently converges to similar solutions with minimal loss, irrespective of the initial values assigned to $w^{l,h}$, $w^{l,r}$, and $\textbf{b}^{l}$. Please refer to appendix B for detailed training information. Additional experiments were conducted to validate the faithfulness of the regression solution further. First, we replaced the MLP outputs $\textbf{m}^{l}$ in each layer with the $\tilde{\textbf{m}}^{l}$, the regression value following Eq 7. We computed the Kullback-Leibler (KL) divergence (Kullback and Leibler, 1951) between the output distributions of the modified model and those of the original model. The resulting average KL divergence is 0.21, indicating a close match between the distributions. Moreover, we analyzed the average final logit of $Y$ produced by both the original and modified models. The logit values are close, measuring 13.08 and 13.15, respectively.
Our primary focus lies on the regression coefficients for layer 9 because MLP9 directly follows the L9H8, the most important argument passer in the country-capital task. The coefficients learned from all test data for layer 9 are reported in Table 1. These coefficients reveal an interesting pattern. Among all the terms, only three exhibit positive values, with $w^{9,8}$ being the highest. This suggests that MLP9 enhances the information sourced from L9H8. Consequently, the arguments passed by L9H8, along with some capital names it generates, “stand out” in the residual stream.
@@
To summarize our progress thus far, we have demonstrated that task-specific heads pass the arguments while other heads also individually write to the residual stream. A subsequent MLP serves as an “activation” for the attention heads’ outputs, effectively highlighting the expected arguments in the residual stream. These findings collectively address the first research question. In the following subsection, we move forward and analyze the layer 10, illustrating how MLP10 applies the “implicit function” via the intercept $\textbf{b}^{l}$.
@@ 3.4 Function Application is “ $\textbf{b}+\textbf{r}_{\textnormal{mid}}$ ”
While partial function application has been identified within the OV circuit of L9H8, the probability of capital cities $Y$ in the decoded residual stream by the end of layer 9 remains lower compared to that of countries, as illustrated in Figure 3(a). It is not until layer 10 that the probability of capital cities $Y$ surpasses that of countries $X$ for the first time, indicating the primary occurrence of function application. In this section, we delve into what happens within layer 10.
@@
We analyzed the cosine similarities among several pairs of vectors, and the results are reported in Figure 4(b). In Figure 4(c), we offer a 3D toy example to grasp the relationships between these vectors. Based on these two figures, we hypothesize how MLP10 accomplishes “function application.” We can see that $\textbf{r}^{10}_{\text{mid}}$ shows a higher degree of similarity with $\textbf{W}_{U}[X]$, a column vector of dimension ${d_{\text{model}}}$, indicating a stronger tendency towards predicting $X$. Despite $\textbf{m}^{10}$ being perpendicular to both $\textbf{W}_{U}[X]$ and $\textbf{W}_{U}[Y]$, its addition to $\textbf{r}^{10}_{\text{mid}}$ could divert the residual stream away from $\textbf{W}_{U}[X]$ and towards $\textbf{W}_{U}[Y]$, resulting in a higher probability of $Y$ in $\textbf{r}^{10}_{\text{post}}$.
@@
Figure 4: (a) The probabilities of $X$ and $Y$ decoded from various vectors. (b) Cosine similarity between the vectors. Note that although the cosine similarity between $\textbf{r}^{10}_{\text{post}}$ and $\textbf{W}_{U}[X]$ remains higher than that with $\textbf{W}_{U}[Y]$, when taking into account the norm, the logit value of $Y$ is higher. (c) A simplified example demonstrating the relationships among vectors in 3D space.
To validate this hypothesis, we project $\textbf{m}^{10}$ onto the hyperplane spanned by $\textbf{W}_{U}[Y]$ and $\textbf{W}_{U}[X]$. Specifically, we stack $\textbf{W}_{U}[Y]$ and $\textbf{W}_{U}[X]$ as columns, constructing the corresponding projection matrix P. The projection of $\textbf{m}^{10}$ onto the hyperplane is achieved by its multiplication with P:
$\textbf{M}=[\textbf{W}_{U}[X],\textbf{W}_{U}[Y]],$ (8)
$\textbf{P}=\textbf{M}(\textbf{M}^{\top}\textbf{M})^{-1}\textbf{M}^{\top},$
$\textbf{m}^{10,\text{proj}}=\textbf{P}\textbf{m}^{10}.$
We should clarify that we do not assert that the intercept invariably directs the residual stream from “$X$” towards “$Y$” during the function application phase. Understanding the underlying mechanism of specific MLP layers calls for individual scrutiny in each case, considering the potentially intricate behaviors that influence the probability of specific tokens. We will discuss this in appendix E.
@@
To sum up, we have addressed the research questions listed earlier.
4 The Universal Anti-Overconfidence Mechanisms at the Final Layer
@@
Readers may notice another phase termed “saturation” in the one-shot scenario explored by Merullo et al. (2023), following “function application,” wherein $Y$ maintains its top rank in the final layers. In contrast, as depicted in Figure 3(a) in our zero-shot scenario, the probabilities of both $X$ and $Y$ drastically drop in the last two layers. We then extend our investigation from the zero-shot to few-shot settings to delve deeper.
4.1 Comparing Zero-shot with Few-shot Scenarios
Zero-shot Mechanisms Still Work in Few-shot Scenarios

### v4

@@ 3.2 Attention heads pass arguments to the implicit function
We have two closely related pieces of evidence suggesting that attention heads, L9H8 and L10H0, play pivotal roles in the two phases of factual recall.
(1) We replicated experiments conducted by Merullo et al. [21] using GPT-2 small. For a given input $p_{\text{original}}$, we assessed the early-decoded logits $\pi(X,\textbf{r}^{l}_{\text{post}})$ and $\pi(Y,\textbf{r}^{l}_{\text{post}})$ at each layer $l$, followed by applying softmax to these logits to obtain the probabilities of $X$ and $Y$ at layer $l$ respectively. Illustrated in Figure 3(a), at layer 9, $X$ becomes noticeable in the residual stream, marking the onset of “argument formation.” At layer 10, the probability of $Y$ surpasses that of $X$ for the first time, indicating the “function application.”
(2) We implement the “path patching” algorithm, an activation patching algorithm introduced by [39] to identify influential nodes. For patching, $p_{\text{original}}$ is sampled from the test dataset, while the set of interventions $\{p_{\text{intervention},n\leq N}\}$ is the entire test dataset. Taking $\Delta\pi$ (Eq. 4) as the impact metric, path patching results illustrated in Figure 3(b) indicate that paths from outputs of L9H8 and L10H0, i.e., $\textbf{a}^{9,8}$ and $\textbf{a}^{10,0}$, to $\textbf{r}^{11}_{\text{post}}$ have the most substantial positive impact on correct prediction.
We identified that L9H8 and L10H0 are “mover heads” [39, 17, 22], which move the information from any position they focus on to the final position. Figure 3(c) shows that both L9H8 and L10H0 predominantly attend to the token $X$; Figure 3(d) illustrates the attention score from the final token to $X$, plotted against the inner product $<\textbf{a}^{l,h}_{t=-1},\textbf{W}_{U}[X]>$. The significant correlation between the attention score and the inner product (with Pearson correlation coefficient $0.77$ for L9H8 and $0.90$ for L10H0) strongly suggests that these two heads are responsible for passing the country name, the “argument” in the country-capital task, to the final position where the current prediction occurs.
@@
While examining L9H8, we decoded $(\textbf{W}^{\text{L9H8}}_{OV}\cdot\textbf{r}^{9}_{\text{pre},t=X})$ by employing $\texttt{Unembed}(\cdot)$, and analyzed 5 tokens with the top logit values. We noticed that correct capital cities emerged among these top-probable tokens. This indicates that partial function application happens within this attention head. We found that mapping countries to capitals is an inherent ability of L9H8: Despite receiving the embedding vector of a country rather than a representation from deeper layers at position $X$, its OV matrix still yields some capital cities. Our investigation further reveals that L9H8 is a task-specific head mainly activated given country-related prompts, as evidenced by its inactivation in other tasks (see Appendix F). In contrast, although L10H0 lacks an inherent ability to map countries to capital cities within its OV matrix, it is activated across a variety of tasks, even those solely reliant on in-context learning [39, 22]. It functions more like an amplifier for L9H8 or, more generally, an amplifier for the confident answer in the residual stream.
@@ 3.3 MLP acts as the “activation” of attention heads
In §3.2, we have demonstrated that only a few attention heads have a positively direct effect on the correct prediction. This naturally leads to a question: Given that Transformer models aggregate the outputs of all attention heads with equal weight (as shown in Eq. 1), how does the model discern and extract information written by specific heads within the residual stream?
We evaluated the norm of outputs from each attention head and applied principal component analysis (PCA) to identify any anomalous outputs but found no meaningful results. Then, we hypothesize an “activation” mechanism after these attention heads that either amplifies or suppresses the outputs of certain attention heads. We turn to the subsequent module following the attention heads—the MLP.
Recall that a Transformer layer’s output $\textbf{r}^{l}_{\text{post}}$ is the linear combination of $\textbf{r}^{l}_{\text{pre}}$, the collective output of attention heads $\sum^{H-1}_{h=0}\textbf{a}^{l,h}$, and the MLP output $\textbf{m}^{l}$ (See Eq. 1). We consider whether $\textbf{m}^{l}$ is equivalent to a linear combination of $\textbf{r}^{l}_{\text{pre}}$ and $\sum^{H-1}_{h=0}\textbf{a}^{l,h}$, where each term’s coefficient indicates the degree of amplification or erasure imposed by the MLP. We assume the linear independence of all $\textbf{a}^{l,h}$ and $\textbf{r}^{l}$, and tackle the following multivariable linear regression problem:
@@
Here, $w^{l,\cdot}$ represents scalar coefficients to be solved, and the intercept $\textbf{b}^{l}\in\mathbb{R}^{d_{\text{model}}}$ is expected to encapsulate higher-order transformations of head outputs, along with abstract information such as knowledge introduced by the MLP. We employ the gradient descent method to solve this linear regression problem, with the objective of minimizing the Mean Squared Error (MSE) loss between $\tilde{\textbf{m}}^{l}$ and $\textbf{m}^{l}$. We set the learning rate to 0.005 and performed 60,000 optimization steps using the SGD optimizer with a momentum of 0.99. A 4-fold cross-validation to validate the regression solution.
The optimization process consistently converges to similar solutions across all layers with minimal loss, regardless of the initialization of $w^{l,h}$, $w^{l,r}$, and $\textbf{b}^{l}$ (refer to Appendix I for training specifics). This consistency suggests a near-uniqueness of solutions, supporting the assumption of linear independence. To further verify the fidelity of the regression solution, additional experiments were conducted. We substituted the MLP outputs $\textbf{m}^{l}$ in each layer with $\tilde{\textbf{m}}^{l}$ calculated using Eq 6. Subsequently, we evaluated the Kullback-Leibler (KL) divergence [16] between the output distributions of this modified model and those of the original model, resulting in a minimal KL divergence of 0.21. Furthermore, the average final logits of $Y$ produced by two models yielded values of 13.08 and 13.15. These results demonstrate the alignment of the two outputs and thus affirm the fidelity of the regression solution.
We primarily focus on the regression coefficients for layer 9 (detailed in Table 1) as MLP9 immediately follows the L9H8, the most crucial argument passer in the country-capital task. Notably, among all coefficients, only three exhibit positive values, with $w^{9,8}$ being the highest. This suggests that MLP9 enhances the information sourced from L9H8. Consequently, the arguments passed by L9H8, along with some capital names it generates, “stand out” in the residual stream.
@@
Thus far, we demonstrated that task-specific heads pass the arguments, and a subsequent MLP serves as an “activation” for the attention outputs, effectively highlighting the expected arguments in the residual stream. These findings collectively address the first research question.
@@ 3.4 Function application is “ $\textbf{b}+\textbf{r}_{\textnormal{mid}}$ ”
While partial function application has been identified within the OV circuit of L9H8, the probability of capital cities $Y$ by the end of layer 9 remains lower compared to that of countries, as illustrated in Figure 3(a). At layer 10, the probability of $Y$ surpasses that of $X$ for the first time, indicating the primary occurrence of function application. In this section, we study what happens within layer 10.
@@
To answer this question, we analyzed the cosine similarities among several pairs of vectors, and the results are reported in Figure 4(b). In Figure 4(c), we offer a 3D toy example to grasp the relationships between these vectors. We hypothesize that despite $\textbf{m}^{10}$ being almost perpendicular to both $\textbf{W}_{U}[X]$ and $\textbf{W}_{U}[Y]$, its addition to $\textbf{r}^{10}_{\text{mid}}$ could divert the residual stream away from $\textbf{W}_{U}[X]$ and towards $\textbf{W}_{U}[Y]$, resulting in a higher probability of $Y$ in $\textbf{r}^{10}_{\text{post}}$.
To validate the above hypothesis, we project $\textbf{m}^{10}$ onto the hyperplane spanned by $\textbf{W}_{U}[Y]$ and $\textbf{W}_{U}[X]$. Specifically, we stack $\textbf{W}_{U}[Y]$ and $\textbf{W}_{U}[X]$ as columns, constructing the corresponding projection matrix P. The projection of $\textbf{m}^{10}$ onto the hyperplane is achieved by its multiplication with P:
$\textbf{M}=[\textbf{W}_{U}[X],\textbf{W}_{U}[Y]],$ (7)
$\textbf{P}=\textbf{M}(\textbf{M}^{\top}\textbf{M})^{-1}\textbf{M}^{\top},$
$\textbf{m}^{10,\text{proj}}=\textbf{P}\textbf{m}^{10}.$
We should clarify that we do not assert that the intercept invariably directs the residual stream from “$X$” towards “$Y$” in deep layers. Understanding the underlying mechanism of specific MLP layers calls for individual scrutiny in each case, considering the potentially intricate behaviors that influence the probability of specific tokens. We will discuss this in Appendix G.
To sum up, we have addressed the research questions listed earlier.
@@
Figure 4: (a) The probabilities of $X$ and $Y$ decoded from various vectors. (b) Cosine similarity between the vectors. Note that although the cosine similarity between $\textbf{r}^{10}_{\text{post}}$ and $\textbf{W}_{U}[X]$ remains higher than that with $\textbf{W}_{U}[Y]$, when taking into account the norm, the logit value of $Y$ is higher. (c) A simplified example illustrating the relationships among vectors in 3D space. As MLP10 generates a component $\textbf{b}^{10,\text{proj}}$ that aligns with $\textbf{W}_{U}[Y]-\textbf{W}_{U}[X]$ in direction, adding $\textbf{b}^{10,\text{proj}}$ to the residual stream achieves the “function application,” causing the probability of $Y$ to surpass that of $X$ for the first time during the forward pass.
4 The universal anti-overconfidence mechanisms at the final layer
Readers may notice another phase termed “saturation” in the one-shot scenario explored by Merullo et al. [21], following “function application,” wherein $Y$ maintains its top rank in the final layers. In contrast, as depicted in Figure 3(a) in the zero-shot scenario, the probabilities of both $X$ and $Y$ drastically drop in the last two layers. To investigate this disparity, we extended our examination from zero-shot to few-shot settings, and uncovered several intriguing findings. Due to page constraints, we briefly introduce these findings and reserve detailed experiments and discussions for Appendix A.
@@ 4 The universal anti-overconfidence mechanisms at the final layer
Zero-shot mechanisms still work in few-shot scenarios. (Appendix A.1)
Figure 5 shows that the identified influential heads remain primarily consistent regardless of the number of demonstrations provided. Additionally, the probability dynamics of $X$ and $Y$ resemble those in zero-shot scenarios: the formation of arguments and function application are clear, and in the final layers, the probabilities of both $X$ and $Y$ decrease. It becomes evident that the identification of the “saturation” phase by Merullo et al. [21] is not due to in-context demonstrations but rather stems from the metric they utilize, the token rank rather than detailed probability. Consequently, the final declines are overlooked.
ICL improves model confidence to generate capital’s name. (Appendix A.2)
We discovered that with more ICL demonstrations, the emergence of probabilities for $X$ and $Y$ occurs earlier, for instance, in layer 8 in few-shot scenarios compared to layer 9 in zero-shot scenarios (see top figures in Figure 5). Meanwhile, the decline in the final layer is relatively slight. We found that in zero-shot scenarios, the model tends to output “safe” tokens such as “not” following “The capital of X is.” In few-shot scenarios, we detected some potential mechanisms that prompt the model to recognize the output domain to be the name of the capital, thereby enhancing the confidence in outputting $Y$.
The final layer avoids overconfidence. (Appendix A.4)
We revealed that the final layer of language models takes two steps to avoid outputting overconfident answers.
Figure 6: The anti-overconfidence mechanism at the last layer imposed by attention heads and MLPs. This mechanism can be suppressed according to our analysis.
(1) Attention heads incorporate frequent tokens into the residual stream. We found that most attention heads in the last layer focus on the first position (see Figure 8) where their OV circuits output frequent English tokens such as "and", "a", and ".", etc. These frequent tokens are copied from the first position to the final position, diluting the probability of $Y$.
(2) The MLP generates a vector that steers the residual stream towards an “average” token, the frequency-weighted average token embedding over the training corpus, denoted as $\mathbb{E}[\textbf{W}_{U}]$. When applying our linear regression analysis to the final MLP layer and obtaining the corresponding intercept term b, there is high similarity between ($\mathbb{E}[\textbf{W}_{U}]$ $\textbf{W}_{U}[Y]$) and b, indicating that this MLP directs the residual stream from the correct answer $Y$ to the “average” token.
These two steps are intuitive: the model incurs a large loss when a confident answer is incorrect. By aligning the model outputs with the expected logits distribution across the training corpus, the model losses could be effectively minimized. They can be suppressed for a confident output by (1) applying attention masks to block information flow from the first position and (2) subtracting b from the residual stream. For example, as shown in Figure 6, on GPT-2 small, we improved the zero-shot output probability of $Y$ from 5.53% to 25.93%.
5 Conclusion
In this paper, we conduct a thorough investigation into the factual recall mechanisms utilized by language models. We studied the mechanisms behind “argument passing” and “function application.” Furthermore, we uncovered the prevalence of anti-overconfidence mechanisms in language models. Our proposed analysis method, based on linear regression, effectively decomposes MLP outputs into components that are easily understandable to humans. This method has been substantiated through numerous empirical experiments and lays a valuable foundation for our interpretations. Finally, we study the anti-overconfidence mechanisms in the final layer, which is a universal mechanism regardless of models, tasks, and the number of in-context demonstrations.
Acknowledgments and Disclosure of Funding
Appendix A The universal anti-overconfidence mechanisms at the final layer
@@
Readers may notice another phase termed “saturation” in the one-shot scenario explored by Merullo et al. [21], following “function application,” wherein $Y$ maintains its top rank in the final layers. In contrast, as depicted in Figure 3(a) in our zero-shot scenario, the probabilities of both $X$ and $Y$ drastically drop in the last two layers. We then extend our investigation from the zero-shot to few-shot settings to delve deeper.
A.1 Zero-shot mechanisms still work in few-shot scenarios